# Weighted Covariance Intersection for Range-based Distributed Cooperative Localization of Multi-Agent Systems

### v1

@@ V-A DCL Analysis
In this subsection, we analyze the performance of the proposed WCI-based DCL algorithm, comparing it with all benchmarks, particularly classical CI. The simulation settings are set to their default values. The RMSE and STD for all state components using different methods are shown in Fig. 5. The averaged RMSE, STD, and NEES over all time steps for all state components using different methods are presented in Table IIIb. The results for agent 1 and agents 2-8 are listed separately due to their different connections to anchors.
Figure 5: The RMSE and STD for all state components using different methods. (a) agent 1. (b) agent 2-8 (averaged).
TABLE III: RMSE, STD and NEES for Different Methods and State Components (Averaged over all time steps)
@@
RMSE STD NEES RMSE STD NEES RMSE STD NEES RMSE STD NEES RMSE STD NEES
NCL 3982 3.0E+04 394.2 117.2 1.0E+04 7.702 30.96 3.0E+05 1.624 2495 1.4E+08 1.480 2425 1.4E+08 1.479
EKF 0.064 0.060 3.495 0.032 0.029 4.569 0.921 0.726 5.481 271.4 266.4 6.410 590.2 456.2 12.01
CI-trace 0.223 1.053 0.296 0.249 1.261 0.243 53.49 169.2 3.643 2.1E+04 6.7E+04 1.879 5.4E+04 1.5E+05 2.178
CI-det 0.458 1.660 0.250 0.120 0.695 0.129 2.692 29.36 0.186 499.6 4672 0.162 1208 1.5E+04 0.150
WCI 0.190 1.289 0.157 0.087 0.636 0.083 2.171 27.27 0.088 441.0 5022 0.080 1023 1.3E+04 0.075
(a) RMSE, STD and NEES of agent 1
Method Position [m] Velocity [m/s] Attitude [deg] Gyroscope Bias [deg/h] Accel Bias $\boldsymbol{\mu}$g]
RMSE STD NEES RMSE STD NEES RMSE STD NEES RMSE STD NEES RMSE STD NEES
NCL 8.394 1.614 8.0E+04 0.330 0.123 1249 1.350 0.909 36.85 312.2 301.7 27.52 962.3 673.3 117.0
EKF 0.053 0.049 3.511 0.028 0.025 4.419 0.946 0.681 5.304 272.4 256.7 6.001 536.5 434.2 8.736
CI-trace 0.144 0.338 2.491 0.124 0.234 4.117 21.19 22.68 8.275 4971 6906 4.922 1.1E+04 1.7E+04 3.322
CI-det 0.568 0.806 2.607 0.066 0.083 2.810 1.157 0.954 3.982 300.2 316.5 3.868 642.8 658.3 4.131
WCI 0.155 0.405 2.002 0.047 0.078 1.860 1.180 1.310 2.148 304.2 377.0 2.058 631.9 881.8 1.741
(b) Averaged RMSE, STD and NEES of agents 2-8
First, compared to the non-cooperative localization, all cooperative localization algorithms significantly improve state estimation, demonstrating that cooperation greatly enhances the usability of state estimation, especially position, under conditions of insufficient measurements with anchors. Among cooperative localization algorithms, it can be observed that the CCL EKF algorithm consistently provides the optimal estimates for all state components. In contrast, DCL algorithms, which do not track inter-agent correlations, show inferior state estimation performance compared to EKF. Furthermore, CI-based algorithms (including WCI) are inherently conservative estimation methods, with STD being higher than RMSE, which is an inherent characteristic of CI. When comparing agent 1 and agents 2-8, the primary difference lies in the availability of measurements from anchors. It is evident that nodes with anchor measurements integrated into the EKF update show some improvement in estimation accuracy, while their estimation conservativeness is significantly reduced. Additionally, when examining the estimation of different state components, it can be observed that the CI-trace method performs well in position estimation, with relatively small RMSE and the smallest STD among DCL algorithms. This is primarily due to the fact that its optimization objective is predominantly driven by position uncertainty. However, for other state components with smaller scales, such as attitude, gyroscope bias, and accelerometer bias, the CI-trace performs poorly. The attitude estimation error reaches several tens of degrees and the bias estimates also fail to serve as an inertial alignment correction. This has a detrimental impact on the agent: On the one hand, many tasks require not only precise position but also accurate attitude control; on the other hand, inaccurate attitude and bias estimates significantly impair the agent’s ability of autonomous navigation. When the agent fails to obtain sufficient distance measurements, the errors by only inertial navigation quickly accumulate and diverge. CI-det, on the other hand, achieves better estimation results for state components with smaller scales including attitude and IMU biases. However, it performs the worst in position estimation. In contrast, WCI achieves a balanced estimation across all states, yielding satisfactory results.
WCI and classical CI differ only in the state update phase. We then conduct an in-depth comparison of their performance during the state update phase. Agent 1 is selected as the subject of study, as its state update involves a single update method, unaffected by the interference of EKF updates based on distance measurements to anchors. Fig. 6 shows the change in the trace of the covariance matrix for each state component after each update, reflecting the variation in the uncertainty of the states. Since the trace changes exhibit significant variability and can be both positive and negative, a symmetric logarithmic scale is employed for the Y-axis (with non-equidistant intervals). We observe that, in contrast to EKF, which consistently reduces the uncertainty of each state component after each update, CI-based methods fail to achieve such consistency, especially for states with smaller scales, such as attitude, gyroscope bias, and accelerometer bias. For these states, it is evident that, in some cases, the updates lead to an increase in uncertainty, with the trace changes displaying an oscillatory pattern. However, compared to CI-trace, both CI-det and WCI better balance the optimization of states with varying scales. States with larger scales, such as position and velocity, show smaller changes in uncertainty after each update, while the oscillations of smaller scale states are reduced. This enables CI-det and WCI to achieve a more effective balance in optimizing uncertainty across different states, as illustrated in Fig. 5 and Table IIIb, thus preventing the severe divergence in uncertainty for smaller magnitude states observed in CI-trace.
Figure 6: The change in the trace of the covariance matrix for each state component after each update
We compared the nominal correlations imposed by the covariance matrices for the CI-trace, CI-det, and WCI methods, along with the actual correlations of the states. The average differences between the nominal and actual correlations are shown in Fig. 7. It is evident that, under all three methods, the actual correlations differ from the nominal ones, which is an inherent issue of CI that is difficult to avoid. However, in this experiment, compared to CI-det, both WCI and CI-trace exhibit lower correlation mismatches, particularly in the correlations related to position (the top three rows of the matrix), which dominate the state update. This explains why CI-det, while reflecting a better optimization balance between states of different scales, shows degradation in position estimation performance. In fact, the actual advantage of WCI lies in its flexibility, allowing the design of a weighting matrix $\boldsymbol{W}$ to balance different scale state estimations and address correlation mismatches in cooperative localization problems.
Figure 7: The average differences between the nominal and actual correlations
@@
We also compared the impact of the hyperparameter $T_{a}$ chosen for the weighting matrix in WCI. The results are shown in Fig. 8, where we evaluate the performance of WCI with different weighting matrix by selecting different values of the hyperparameter $T_{a}$ in (30) and (32). Since CI-based methods are conservative estimators, STD and RMSE do not perfectly correspond to each other. WCI primarily optimizes based on STD. A larger value of $T_{a}$ indicates a greater emphasis on the long-term performance of the inertial navigation system during optimization, which places more weight on the uncertainty of states with higher-order relationships to propagation errors, such as attitude and IMU biases. This is reflected in the figure, where, as $T_{a}$ increases, the uncertainty in the position estimate gradually increases, while the uncertainty in the estimates of attitude and IMU biases decreases.
Figure 8: State estimation performance versus different values of $T_{a}$ for WCI
@@
At the end of this subsection, we evaluate the advantages of the concurrent update in the proposed DCL algorithm. We compare it with the sequential update counterpart, where each agent updates its states immediately upon receiving one measurement. As shown in Fig. 9, EKF demonstrates little difference between the two schemes. However, for WCI, it is evident that the proposed concurrent update better aligns with the characteristics of CI fusion, exhibiting lower conservatism and superior estimation performance compared to sequential updates.
@@ V-B1 Robustness
Figure 10: Position error comparison between DCL and CCL algorithms. After 30 seconds of motion, there is a 5% probability of data transmission failure from agent 1 to the data center. At 90 seconds, agent 1 goes offline and loses connection with both the computing center and other agents. It then rejoins the network at 110 seconds, restoring its connection with other agents.
The CCL algorithm requires all information to be gathered at the computing center. If packet loss or certain agent failure occurs, it not only halts the state estimation of the affected agent but also leads to incorrect correlation maintenance between agents. This, in turn, gradually affects the state estimation of other agents as well. We evaluate the robustness of the distributed and centralized algorithms under two conditions: (1) Communication packet loss, where data from certain nodes fails to reach the computing center, and (2) Changes in the network topology, where nodes go offline or join the system. It is assumed that, starting from 30 seconds, there is a 5% probability of data transmission failure from agent 1 to the data center. At 90 seconds, agent 1 goes offline, losing its connection with the computing center and other agents. Agent 1 then re-joins the network at 110 seconds, restoring its previous connections.
As shown in Fig. 10, we observe that after 30 seconds, not only does the position estimation error of agent 1 gradually increase, but it also skews the estimation results of other agents that have not experienced packet loss. At this point, the actual RMSE exceeds the estimated STD, indicating inconsistent estimates. In contrast, the DCL methods are completely unaffected by this issue.
During the 90-110s period, agent 1 is fully offline and disconnected with all anchors and other agents, relying solely on its own IMU outputs for state propagation. The state estimation error increases rapidly. However, methods like WCI and CI-det, which achieved balanced state estimation in earlier stages, show smaller attitude and IMU biases errors and uncertainties, leading to slower error accumulation during the autonomous navigation phase. Furthermore, when agent 1 reconnects, its state is not updated by the centralized algorithm during its offline period. As a result, when agent 1 re-joins, its differential distance measurements with other agent exhibit significant bias, which in turn skews the position estimates of the other agents. In contrast, the DCL algorithm allows the estimation of each agent to remain independent, making it more robust.
@@ V-B2 Computation and Communication Overhead
To evaluate computational complexity, we compared the total processing time of the DCL algorithm and the CCL algorithm under different swarm sizes, as shown in Fig. 11. The total computational cost of the DCL algorithm increases linearly with the number of agents. When averaged on each agent, the computational overhead for its own state estimation remains almost constant as the swarm size increases, making it more suitable for large-scale swarms. In contrast, as the number of agents increases, the dimension of the covariance matrix in the CCL algorithm grows linearly, resulting in a cubic growth in computational cost (since the computational complexity of matrix multiplication and inversion involved in state prediction and update are all cubic in terms of matrix dimensions [32]). As a result, the processing time grows rapidly, making it difficult to scale to large swarms.
As for the communication load, the CCL algorithm requires high-frequency IMU data and measurement data to be collected at the computing center. Additionally, agents rely on messages from the computing center to obtain their state estimates. Some distributed implementations of CCL algorithms can also broadcast their state and state transition matrices to track the correlations between agents [33], which, although optimized to some extent, still require the transmission of large amounts of data. In contrast, our distributed approach only requires each agent to broadcast its position and attitude estimates, along with their variances, amounting to just a few dozen bytes of data, significantly reducing the communication burden of the system. Since this conclusion is evident, no separate simulation experiments are conducted for further analysis.
@@ VI Conclusion
In this paper, we propose a WCI-based DCL algorithm designed for MAS operating in 3D scenarios. We identify three major drawbacks of CI in recursive filtering, which may pose risks to recursive filtering tasks, such as poor performance in 3D cooperative localization. We discover that a weighted mechanism, specifically WCI, has the potential to mitigate the inherent drawbacks of CI in recursive filtering tasks. Building on this, we introduce WCI into range-based DCL in 3D scenarios, designing a fusion strategy for multiple measurements and a weighting matrix selection strategy guided by the error propagation rule of INS for the cooperative localization task. Simulation results confirm that the proposed WCI algorithm outperforms classical CI in 3D cooperative localization tasks, and the proposed distributed framework achieves higher precision and tighter uncertainty compared to sequential updates. Additionally, the distributed framework demonstrates greater robustness and scalability, making it suitable for large-scale MAS. Future work will focus on developing dynamic weighting matrix selection strategies to adapt to the agents’ motion and cluster configuration changes, as well as applying more advanced distributed data fusion methods to improve the conservative nature of CI.
Appendix A Proof of Proposition 2
Reinhardt et al. studied the optimality of CI in [34]. We use its conclusion as a lemma:
Lemma 1.
For any strictly monotonically increasing cost function $J(\cdot)$ ($\boldsymbol{P}_{1}\succ\boldsymbol{P}_{2}\Rightarrow J(\boldsymbol{P}_{1})>J(\boldsymbol{P}_{2})$), CI provides the optimal bound in the fusion of two estimates under unknown correlations.
Based on Lemma 1, to prove Proposition 2, we should only prove that for any positive-definite weight matrix $\boldsymbol{W}$, $\operatorname{tr}(\boldsymbol{W}\boldsymbol{P})$ is a strictly monotonically increasing cost function, i.e., if and only if $\boldsymbol{W}$ is positive-definite, for any positive-definite matrices $\boldsymbol{P}_{1}\succ\boldsymbol{P}_{2}$, it holds that
$\operatorname{tr}(\boldsymbol{W}\boldsymbol{P}_{1})>\operatorname{tr}(\boldsymbol{W}\boldsymbol{P}_{2})\leftrightarrow\operatorname{tr}(\boldsymbol{W}(\boldsymbol{P}_{1}-\boldsymbol{P}_{2}))>0$ (37)
We prove Proposition 2 in two parts: sufficiency and necessity.
  • Sufficiency
Since $\boldsymbol{P}_{1}\succ\boldsymbol{P}_{2}$, $\boldsymbol{\Delta}=\boldsymbol{P}_{1}-\boldsymbol{P}_{2}$ is positive-definite. The positive-definite matrix $\boldsymbol{W}$ can be expressed as $\boldsymbol{W}=\boldsymbol{Q}^{\mathsf{T}}\boldsymbol{Q}$ for certain invertible $\boldsymbol{Q}$. Then
$\begin{split}\operatorname{tr}\left(\boldsymbol{W}(\boldsymbol{P}_{1}-\boldsymbol{P}_{2})\right)&=\operatorname{tr}\left(\boldsymbol{Q}^{\mathsf{T}}\boldsymbol{Q}\boldsymbol{\Delta}\right)\\ &=\operatorname{tr}\left(\boldsymbol{Q}\boldsymbol{\Delta}\boldsymbol{Q}^{\mathsf{T}}\right)\end{split}$ (38)
The second equation applies the commutative law of matrix trace [32]. Let $\boldsymbol{S}=\boldsymbol{Q}\boldsymbol{\Delta}\boldsymbol{Q}^{\mathsf{T}}$. Since $\boldsymbol{\Delta}$ is positive-definite and $\boldsymbol{Q}$ is invertible, $\boldsymbol{S}$ is positive-definite as well. Therefore, $\operatorname{tr}(\boldsymbol{S})=\operatorname{tr}(\boldsymbol{W}(\boldsymbol{P}_{1}-\boldsymbol{P}_{2}))>0$.
  • Necessity
If the weight matrix $\boldsymbol{W}$ is not positive-definite, then there must exist $\boldsymbol{x}\neq\boldsymbol{0}$ which satisfies $\boldsymbol{x}^{\mathsf{T}}\boldsymbol{W}\boldsymbol{x}\leq 0$. Let $\boldsymbol{\Delta}=\boldsymbol{P}_{1}-\boldsymbol{P}_{2}=\boldsymbol{x}\boldsymbol{x}^{\mathsf{T}}$ ($\boldsymbol{\Delta}$ is clearly positive-definite), then
$\operatorname{tr}(\boldsymbol{W}\boldsymbol{\Delta})=\operatorname{tr}(\boldsymbol{W}\boldsymbol{x}\boldsymbol{x}^{\mathsf{T}})=\boldsymbol{x}^{\mathsf{T}}\boldsymbol{W}\boldsymbol{x}\leq 0$ (39)
which is conflicted with the condition that $\operatorname{tr}(\boldsymbol{W}(\boldsymbol{P}_{1}-\boldsymbol{P}_{2}))>0$ for any positive-definite matrices $\boldsymbol{P}_{1}\succ\boldsymbol{P}_{2}$. So the assumption is not valid, and the weight matrix $\boldsymbol{W}$ must be positive-definite.

### v2

@@ V-A DCL Analysis
In this subsection, we analyze the performance of the proposed WCI-based DCL algorithm, comparing it with all benchmarks, with a particular focus on traditional CI. The simulation parameters are set to their default values. Due to different connectivity with anchors, we select three representative vehicles, indexed as 1, 3, and 6, for analysis. The RMSE and RMTE for all state components using different methods are shown in Fig. 6. The averaged RMSE, RMTE, and NEES over all time steps for all state components using different methods are presented in Table IIIc.
Figure 6: The RMSE and RMTE for all state components using different methods. (a) vehicle 1. (b) vehicle 3. (c) vehicle 6
First, compared to the non-cooperative localization, all cooperative localization algorithms significantly improve state estimation, demonstrating that cooperation greatly enhances the usability of state estimation, especially position, under conditions of insufficient measurements with anchors. Among cooperative localization algorithms, it can be observed that the CCL EKF algorithm consistently provides the optimal estimates for all state components. In contrast, DCL algorithms, which do not track inter-vehicle correlations, show inferior state estimation performance compared to CCL EKF. Furthermore, CI-based algorithms (including WCI) are inherently conservative estimation methods, with RMTE being higher than RMSE, which is an inherent characteristic of CI. When comparing vehicle 1 and vehicles 3,6, the primary difference lies in the availability of measurements from anchors. It is evident that nodes with anchor measurements integrated into the EKF update generally show some improvement in estimation accuracy, while their estimation conservativeness is significantly reduced. Additionally, when examining the estimation of different state components, it can be observed that the CI-trace method performs well in position estimation, with relatively small RMSE and the smallest RMTE among DCL algorithms. This is primarily due to the fact that its optimization objective is predominantly driven by position uncertainty. However, for other state components with smaller scales, such as attitude, gyroscope bias, and accelerometer bias, CI-trace performs poorly. The attitude estimation error reaches several tens of degrees and the bias estimates also fail to serve as an inertial alignment correction. This has a detrimental impact on the vehicle: On the one hand, many tasks require not only precise position but also accurate attitude control; on the other hand, inaccurate attitude and bias estimates significantly impair the vehicle’s ability of autonomous navigation. When the vehicle fails to obtain sufficient distance measurements, errors from inertial navigation only quickly accumulate and diverge. CI-det, on the other hand, achieves better estimation results for state components with smaller scales including attitude and IMU biases. However, it performs the worst in position estimation, especially for vehicle 3, where its position RMSE is significantly higher than that of other methods. The difference between vehicle 3 and vehicle 6 also indicates the instability of CI-det in position estimation. In contrast, DCL WCI achieves a balanced and stable estimation across all state components, yielding satisfactory results.
TABLE III: RMSE, RMTE and NEES for Different Methods and State Components (Averaged over all time steps)
@@
RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES
 NCL 4750 3886 25.172 138.0 113.8 56.29 30.97 27.94 4.128 2494 2494 3.000 2424 2425 2.998
CCL EKF 0.078 0.071 3.902 0.035 0.032 5.139 1.020 0.758 5.945 290.2 274.5 5.364 599.6 461.8 11.45
DCL CI-trace 0.305 1.419 0.282 0.369 1.923 0.203 73.876 216.4 3.495 3.4E+04 1.1E+05 1.631 7.6E+04 2.3E+05 1.688
DCL CI-det 0.636 2.305 0.319 0.146 0.894 0.131 2.930 32.72 0.186 521.7 4342 0.183 1307 1.6E+04 0.175
DCL WCI 0.259 1.772 0.196 0.108 0.909 0.084 2.854 41.53 0.081 559.6 7063 0.081 1257 1.8E+04 0.076
(a) RMSE, RMTE and NEES of vehicle 1
RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES
 NCL 5.837 2.316 107.5 0.230 0.146 11.20 1.286 0.934 6.389 317.0 306.1 7.180 874.6 712.3 10.38
CCL EKF 0.074 0.062 3.862 0.028 0.026 4.898 0.886 0.695 5.094 258.4 260.5 4.918 550.6 431.3 9.634
DCL CI-trace 0.217 0.632 2.771 0.190 0.464 2.549 46.88 40.91 26.30 1.7E+04 1.7E+04 7.584 3.2E+04 3.2E+04 20.84
DCL CI-det 2.301 1.915 6.238 0.124 0.129 3.469 1.072 0.947 3.912 297.6 310.1 4.427 743.1 705.6 6.028
DCL WCI 0.201 0.727 1.620 0.053 0.113 1.339 1.187 1.636 1.241 303.0 409.8 1.281 682.4 1116 0.882
(b) RMSE, RMTE and NEES of vehicles 3
RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES
 NCL 1.266 1.040 4.584 0.131 0.123 4.035 1.354 0.991 4.821 364.9 317.0 5.553 769.3 729.5 5.502
CCL EKF 0.063 0.059 3.428 0.029 0.026 4.307 0.935 0.680 4.861 270.1 256.3 5.502 498.1 445.0 6.651
DCL CI-trace 0.260 0.481 2.207 0.194 0.341 2.983 26.67 30.39 32.32 8217 8524 5.647 1.3E+04 2.4E+04 4.188
DCL CI-det 0.384 0.556 2.412 0.075 0.094 2.370 1.273 1.075 3.181 347.4 343.6 3.220 660.7 749.6 2.627
DCL WCI 0.203 0.464 1.702 0.066 0.106 1.383 1.394 1.675 1.555 364.7 446.1 1.349 695.3 1144 0.895
(c) RMSE, RMTE and NEES of vehicles 6
To vividly compare the differences in state estimation capabilities between various methods, we use the reciprocal of the averaged RMSE of each state component as a measure of capability. Since the CCL EKF method provides the most accurate state estimation and, therefore, the highest capability, we normalize the capability values using the centralized EKF method as the reference. The resulting radar chart is shown in Fig. 7. From this chart, it is evident that, compared to DCL methods based on traditional CI, DCL WCI achieves a more balanced and effective estimation for all state components, addressing the performance shortcomings of traditional methods in estimating certain states.
Figure 7: The radar chart for state estimation capability of different methods. The capability values of DCL WCI almost entirely encompass those of traditional CI across all five dimensions, effectively addressing the significant shortcomings of the latter. This demonstrates the effective balance that DCL WCI achieves in state estimation.
@@
We also compared the impact of the hyperparameter $T_{a}$ chosen for the weighting matrix in DCL WCI. The results are shown in Fig. 8, where we evaluate the performance of DCL WCI with different weighting matrix by selecting different values of the hyperparameter $T_{a}$ in (30) and (32). Since CI-based methods are conservative estimators, RMTE and RMSE do not perfectly align with each other, and WCI primarily optimizes RMTE. A larger value of $T_{a}$ indicates a greater emphasis on the long-term performance of the inertial navigation system during optimization, thereby assigning more weight to the uncertainty of states with higher-order relationships to propagation errors, such as attitude and IMU biases. This is reflected in the figure, where, as $T_{a}$ increases, the RMTE of position gradually increases, while the RMTEs of attitude and IMU biases decrease.
Figure 8: State estimation performance versus different values of $T_{a}$ for DCL WCI
@@
At the end of this subsection, we evaluate the advantages of the concurrent update strategy in the proposed DCL algorithm. We compare it with the sequential update counterpart, where each vehicle updates its states immediately upon receiving one measurement. As shown in Fig. 9, it is evident that the proposed concurrent update aligns better with the characteristics of CI fusion, exhibiting lower conservativeness and superior estimation performance compared to sequential updates.
@@ V-B1 Robustness
Figure 10: Position error comparison between DCL and CCL algorithms. After 30 seconds of motion, there is a 5% probability of data transmission failure from vehicle 1 to the data center. At 90 seconds, vehicle 1 goes offline and loses connection with both the computing center and other vehicles. It then rejoins the network at 110 seconds, restoring its connection with other vehicles.
The CCL algorithm requires all information to be gathered at the computing center. If packet loss or certain vehicle failure occurs, it not only halts the state estimation of the affected vehicle but also leads to incorrect correlation maintenance between vehicles. This, in turn, gradually affects the state estimation of other vehicles as well. We evaluate the robustness of the distributed and centralized algorithms under two conditions: (1) Communication packet loss, where data from certain nodes fails to reach the computing center, and (2) Changes in the network topology, where nodes go offline or join the system. It is assumed that, starting from 30 seconds, there is a 5% probability of data transmission failure from vehicle 1 to the data center. At 90 seconds, vehicle 1 goes offline, losing its connection with the computing center and other vehicles. vehicle 1 then re-joins the network at 110 seconds, restoring its previous connections.
We select vehicle 1 and vehicle 3 as representative examples for analysis. As shown in Fig. 10, we observe that after 30 seconds, not only does the position estimation error of vehicle 1 by CCL EKF method gradually increase, but it also skews the estimation results of other vehicles (such as vehicle 3) that have not experienced packet loss. At this point, the actual RMSE exceeds the estimated RMTE, indicating inconsistent estimates. In contrast, the DCL methods are completely unaffected by this issue.
During the 90-110s period, vehicle 1 is fully offline and disconnected with all anchors and other vehicles, relying solely on its own IMU outputs for state propagation. The position error increases rapidly. However, methods like DCL WCI and DCL CI-det, which achieved balanced state estimation in earlier stages, preserve smaller attitude and IMU biases errors and uncertainties, leading to slower error accumulation during the autonomous navigation phase. Furthermore, when vehicle 1 reconnects, its state is not updated by the centralized algorithm during its offline period. As a result, when vehicle 1 re-joins, its differential distance measurements with other vehicle exhibit significant bias, which in turn skews the position estimates of the other vehicles (such as vehicle 3). In contrast, the DCL algorithm allows the estimation of each vehicle to remain independent, making it more robust.
@@ V-B2 Computation and Communication Overhead
To evaluate computational complexity, we compared the total processing time of the DCL algorithm and the CCL algorithm under different swarm sizes using the same platform (MATLAB R2022b, Intel Core i7-12700, 32GB RAM), as shown in Fig. 11. The total computational cost of the DCL algorithm increases linearly with the number of vehicles. When averaged on each vehicle, the computational overhead for its own state estimation remains almost constant as the swarm size increases, making it more suitable for large-scale swarms. In contrast, as the number of vehicles increases, the dimension of the covariance matrix in the CCL algorithm grows linearly, resulting in a cubic growth in computational cost (since the computational complexity of matrix multiplication and inversion involved in state prediction and update are all cubic in terms of matrix dimensions [35]). As a result, the processing time grows rapidly, making it difficult to scale to large swarms.
As for the communication load, the CCL algorithm requires high-frequency IMU data and measurement data to be collected at the computing center. Additionally, vehicles rely on messages from the computing center to obtain their state estimates. Some distributed implementations of CCL algorithms can also broadcast their state and state transition matrices to track the correlations between vehicles [36], which, although optimized to some extent, still require the transmission of large amounts of data. In contrast, our distributed approach only requires each vehicle to broadcast its position and attitude estimates, along with their variances, amounting to just a few dozen bytes of data, significantly reducing the communication burden of the system. Since this conclusion is evident, no separate simulation experiments are conducted for further analysis.
@@ VI Conclusion
In this paper, we propose a WCI-based DCL algorithm designed for MVS operating in 3D scenarios. Unlike existing studies that treat CI merely as a plug-in fusion method, we focus on advancing the CI-based data fusion process for 3D cooperative localization. We find that traditional CI methods fail to achieve satisfactory performance in 3D cooperative localization due to significant scale and observability disparities among state components. To address this, we introduce a weighting mechanism, specifically WCI, to range-based DCL in 3D scenarios. Specifically, we design a multi-measurement fusion strategy based on the characteristic of CI fusion and a weighting matrix selection strategy guided by the INS error propagation principles. Simulation results confirm that the proposed DCL WCI algorithm outperforms DCL methods based on traditional CI in 3D cooperative localization tasks, achieving balanced and improved state estimation across different state components. The proposed concurrent fusion strategy yields higher precision and tighter uncertainty bounds compared to sequential updates. Additionally, the distributed framework demonstrates greater robustness and scalability, making it well-suited for large-scale MVS. Future work will focus on developing dynamic weighting matrix selection strategies to adapt to changes in vehicle motion and cluster configurations, as well as exploring more advanced distributed data fusion methods to improve the conservativeness of CI.